solve 5x+5y=1 and -40x-40y=-7 using the addition method
Answer by LinnW(1048)

solve 5x+5y=1 and -40x-40y=-7 using the addition method
multiply 5x+5y=1 by 8
40x+40y=8
add this to -40x-40y=-7
*******************
40x+40y=8
-40x-40y=-7
0=1 So we have an inconsistent set of equations